HETAV PATEL

Grade 11,

16. 12 defective pens are accidentally mixed with 132 good ones. It is not possible to just look at a pen and tell whether or not it is defective. One pen is taken out at random from this lot. Determine the probability that the pen taken out is a good one.

16. 12 defective pens are accidentally mixed with 132 good ones. It is not possible to just look at a pen and tell whether or not it is defective. One pen is taken out at random from this lot. Determine the probability that the pen taken out is a good one.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Numbers of pens = Numbers of defective pens + Numbers of good pens ∴ Total number of pens = 132+12 = 144 pens P(E) = (Number of favourable outcomes/ Total number of outcomes) P(picking a good pen) = 132/144 = 11/12 = 0.916